A special deck of cards has 10 green cards , 12 blue cards , and 5 red cards . When a card is picked, the color is recorded. An experiment consists of first picking a card and then tossing a coin.

a. How many elements are there in the sample space?

b. Let A be the event that a green card is picked first, followed by landing a head on the coin toss. P(A) = Round your answer to 4 decimal places.

c. Let B be the event that a red or blue is picked, followed by landing a head on the coin toss.

Are the events A and B mutually exclusive?

Yes, they are Mutually Exclusive

No, they are not Mutually Exclusive

d. Let C be the event that a green or blue is picked, followed by landing a head on the coin toss.

Are the events A and C mutually exclusive?

Yes, they are Mutually Exclusive

No, they are not Mutually Exclusive
Answer #1

a)number of elements =N(3 elements of color)*N(2 elements of coin toss)=3*2 =6

b)P(A)=(10/27)*(1/2)=5/27

c) Yes, they are Mutually Exclusive

d) No, they are not Mutually Exclusive